Question 1 (Total 15 marks) Financial statement for White Eagle Limited at 2017 fiscal year end as following: White Eagle Limited Balance Sheet as of December 31, 2017 (US dollar in millions) Cash S 1,40 L Accounts payable 1,902 Receivables 1,682C Inventories 1,673 Other current liabilities 298 Total CA4,762 Net fixed assets 27,595 Total assets 32,357 Total CL 2,200 12,715 AseT Long-term debt Common equity17,442 Total liabilities and 32,357 equity White Eagle Limited Income Statement for Year Ended December 31, 2016 (US dollar in millions) Sales Cost of goods sold Gross proft Other operation expenses $ 9,300 (6,904) $ 2,396 (671) $ 1725 (354) $1371 (587) $784 EBIT Interest expense EBT Tax expense Net Income b. Calculate debt ratiog TIE and ROE. The manager of the company is considering issue new debt to repurchase 10% of the equity, assume that sales, operating cost, total assets and tax rate would not affected, the interest payment increases byT2% y how much would the ROE change in response to the change in the eapital structure? How about ROA? (5 marks)
Answer #1

Total Debt = Total Long term liabilities + Current Liabilities = 12,715 + 2,200 = $ 14,915

Debt Ratio = Total Debt / Total Assets = 14915 / 32357 = 0.460951= 46.1 %

TIE = EBIT/ Interest = 1725 / 354 = 4.87

ROE = Net Income / Equity = 784 / 17442 = 0.044949 = 4.49%

When Equity is reduced by 10%, new equity will be = Equity * 90% = 17442 * 0.9 = $15,697.8

New Interest Payment = Interest *1.12 = 354*1.12 = $396.48

Tax rate =Tax expense / EBT = 587 / 1371 = 42.8%

New Income Statement is as follows

Income Statement
Sales 9300
COGS -6904
Gross Profit 2396
Other Operating Expense -671
EBIT 1725
Interest Expense -396.48
EBT 1328.52
Tax 568.6 (42.8%)
Net Income 759.91

New ROE = new Net Income / new Equity =759.91/ 15697.8 =0.0484 = 4.84%

Change in ROE = (New ROE - Original ROE) / Original ROE= (4.84-4.49)/4.49 = 0.078 =7.8%

Thus ROE increases changes by 7.8%

New ROA = Net Income / Total Assets = 759.91/ 32357 = 2.34%